# NeuroDeX: Unlocking Diverse Support in Decompiling Deep Neural Network Executables

### v1

@@ III-D Model Reconstruction
For the global_scale method, different operators have same quant_scale, so the updated weights are close to the original model’s weights. For the kl_divergence method, quant_scale adjusts for different operators. Due to compilation optimizations, different quant_scale can be confused with regular multiplications, making it difficult to identify them. Eventually, the updated weights are a multiple of the actual weights, and different layers have different multiples. To address this, NeuroDeX adopts a substitute training method illustrated in Figure 3. The target model is the quantized compiled DNN executable, NeuroDeX first prepares query dataset to collect labels in the target executable. Next, NeuroDeX freezes known weights and trains the multiples between updated weights and actual weights.
Typical black-box substitute method for model extraction (a in Figure 3) requires retraining all the model weights, whereas NeuroDeX only needs to train a limited multiples (b in Figure 3). A small amount of training data is enough to gain a model that effectively replicates the functionality of the target model.
@@ V-A RQ1: Correctness
BTD uses machine learning to perform operator type recognition and employs symbolic execution methods for operator attribute recovery. BTD can successfully reconstruct nearly identical high-level models from executables compiled by TVM and GLOW in x86 architecture.
However, BTD has some inherent shortcomings. In the operator type recognition stage, BTD trains machine learning model for each compiler version to make predictions. This approach relies heavily on training data and treats the compiler version as prior knowledge, which limits its scalability in real world scenario. Moreover, we find that about 57.9% of the training data in TVM and 18.3% in GLOW appear in the test dataset, further undermining the effectiveness of the method. To avoid the influence of data leak, we randomly split all the data with an 8:2 ratio for train dataset and test dataset, and retrain the model for each compiler version, strictly following the default settings of BTD. As illustrated in Figure 4, we have analyzed BTD’s TRA across different compiler versions. It is evident that BTD’s method exhibits poor cross-version support for compilers after avoiding data leak. Even within the same version, errors of operator type occur to an unacceptable degree, make it difficult to apply directly.
To make the comparison with NeuroDeX more intuitive, we remove the training data that is also on test dataset (operators of the six models in Table IV) and retrain the BTD operator type recognition models following the default settings. The comparison results of TRA are shown in Table IV. The TRA of NeuroDeX significantly outperforms BTD, providing a crucial foundation for accurate operator type recognition, which is essential for subsequent operator attribute recovery and model reconstruction. NeuroDeX achieves results very close to 100% across all experimental models, whereas the BTD method faces relatively severe errors.
@@ V-B RQ2: Efficiency
We choose four models: EfficientNet, ResNet18, Inceptionv1 and ShuffleNetv2. These models cover a range of weights size and topological complexities, enabling a comprehensive evaluation of the overhead associated with BTD and NeuroDeX. The model reconstruction strategies for NeuroDeX and BTD are identical. Therefore, we only compare the time associated with operator type recognition and operator attribute recovery processes. NeuroDeX involves interaction with LLM, which leads to non-fixed time consumption. We measure the time overhead by recording it three times and calculating the average to ensure a fair comparison. BTD uses IDA [20] to decompile the executables and NeuroDeX uses Ghidra. We overlook the difference from general decompiler preprocessing for fair comparison. The compiler version for our experiment is TVM v0.9dev and GLOW 2022. The overhead of BTD and NeuroDeX is shown in Table V.
TABLE V: Overhead Analysis on BTD and NeuroDeX (all value in s)
@@ V-C RQ3: Comprehensiveness
We first evaluate NeuroDeX on the latest compiler verison to verify its scalability. According to our observations, TVM is a project that is frequently maintained; GLOW is a stable project, we have counted the commits since 2023, which total only about 100, and the majority are related to feature maintenance and bug fixes. The repository of GLOW has been marked as public archive now. Therefore, we also use NeuroDeX to analyze the latest TVM version (v0.17). We utilize NeuroDeX to decompile RQ1’s six models in TVM 0.17 repeatedly. We evaluate NeuroDeX on six more different models in TVM v0.17, aiming to cover more model structures and varying hyperparameters for same model structure. The results are shown in Table VI.
TABLE VI: Evaluation of NeuroDeX on the Recent Compiler Version (all value in %)
@@
TABLE VII: Evaluation of NeuroDeX on AArch64 Models (all value in %)
@@
In order to show the cross-architecture support of NeuroDeX, we also evaluate NeuroDeX on aarch64 in TVM v0.17, the results are shown in Table VII. Based on the experimental results, NeuroDeX can successfully decompile executables across different optimization levels on aarch64 architecture. Some errors may accur in operator type recognition stage. The errors include the incorrect identification of Avgpool as Sum, as well as misjudgments of activation functions.
@@
For global_scale method, NeuroDeX can decompile all the four executables with MIA_top1 at least 26.4%, and MIA_top5 at least 40.6%. For kl_divergence method, NeuroDeX can decompile all the four executables with MIA_top1 at least 73.0%, and MIA_top5 at least 96.4%. Given that all models have 1000 output categories, the recovery results are quite satisfactory. In contrast, using the same computational graph with target model and random training with 5000 images (same query dataset with NeuroDeX), the recovered model achieves a maximum MIA_top1 of only 0.6% and a maximum MIA_top5 of only 4.4%, indicating that it almost fails to recover any valuable information. For ResNet18 and ResNet34, NeuroDeX can recover quite similar high-level models under both global_scale and kl_divergence methods. For MnasNet and MobileNetv2, the accuracy of recovered models decreases, we analyze the reasons as follows: compared to ResNet18 and ResNet34, MnasNet and MobileNetv2 have fewer parameters, which makes them more sensitive to the global_scale method. The global_scale method tends to weaken the decision boundaries of models, the queries for recovered model are more likely to be classified into different categories than original high-level models.
@@
V-D RQ4: Robustness
In operator type recognition, NeuroDeX may occasionally encounter errors. Additionally, due to the inherent output variability of LLMs, repeated analysis might yield different error samples. Nonetheless, our evaluation indicates that the operator type recognition accuracy on all TVM operators reach 99.22%, and it is 97.62% for GLOW, maintaining high accuracy. In operator attribute recovery, the errors are very rare and they are obvious in the decompiled code, such as kernel_size “*0.00591716 (1/169)” should be 13 but is incorrectly identified as 17. Therefore, it does not require a systematic error fix mechanism. We analyze the causes of errors in operator type recognition and the errors can be classified into three categories.
  1. Type 1.
    A single operator is split into multiple functions.
  2. Type 2.
    Incorrect recognition of activation functions.
  3. Type 3.
    Other random operator recognition errors.
Type 1: Error operators exhibit fixed patterns. These errors accur in Softmax, Avgpool and Dense_add. These errors can be detected immediately after operator type recognition. NeuroDeX can fix them by matching predefined patterns of operator partition.
To demonstrate the effectiveness of NeuroDeX on different LLMs, we also compare the performance of different LLMs including GPT-4.1 [11], Deepseekv3 [24], GPT-4o mini [13], Gemini2.5 flash [8]. The selection of LLMs will affect TRA directly. Results are shown in Table VIII. Deepseekv3 and GPT-4.1 perform well in all models across different compiler settings. Gemini2.5 flash struggles to identify operator of GLOW, GPT-4o mini fails to identify TVM optimization operators and GLOW operators. In summary, NeuroDeX does not rely on a specific LLM. LLMs that perform well in general domains are equally suitable for completing the operator type recognition task.
TABLE VIII: TRA of Different LLMs (all value in %)
@@ VI-A Possible defenses towards NeuroDeX
NeuroDeX targets standard executables compiled using DL compilers and is capable of recovering nearly identical high-level models. Techniques such as binary obfuscation significantly reduce the readability of decompiled code by general-purpose decompilers like Ghidra. This can make NeuroDeX’s operator type recognition and operator attribute recovery components ineffective. However, DL compilers do not offer direct support for executable obfuscation, and methods like control flow obfuscation may also negatively impact the speed and accuracy of model inference. Overall, NeuroDeX is designed to handle the most common scenarios effectively. The threat model used in our study is consistent with previous works [42, 27, 53, 54] and is generally common and practical in real-world scenarios.
@@ VI-B Scalability of NLP models
Related studies [46, 5] indicate that CV models are the majority models on edge devices, and our evaluation focuses on CV models. In contrast, language models have larger search spaces of input and output, as well as greater parameters. Transformer-based models feature complex operator topologies, and operations like attention are decomposed into several low-level operators, further increasing the complexity of the models. This complexity makes it exceedingly difficult to precisely reconstruct an identical model. Currently, NeuroDeX cannot cover this scenario. However, NeuroDeX’s core components still provide valuable insights, NeuroDeX can enhance the efficacy of gray-box attacks towards these models. We will address this in future work.
As for LLMs with a large number of parameters, to the best of our knowledge, there is currently no mature solution for compiling them into standalone executables in binary format for inference on edge devices. Existing inference engines of LLMs primarily focus on optimizing the computation process but model weights are typically stored and loaded separately.
VI-C Downstream Tasks
NeuroDeX can reconstruct high-level model almost identical to the original model, meaning that computational graph, gradient, weights have been fully exposed, which means that attackers may conduct white-box attacks towards DNN executables on edge devices. Model extraction attacks violate the intellectual property of the model owner. Adversarial examples generated for white-box models can be directly used as inputs for DNN executables, misleading the model’s inference results. Moreover, a white-box model allows for weight modification via backdoor injection. Attackers can edit the executables at the addresses where those weights are stored to inject the backdoor. Backdoor injection results in malicious inputs to manipulate the model’s behavior, compromising its integrity and reliability. Based on this analysis, NeuroDeX reveals the potential security threats posed to DL compilers. Our discussion of NeuroDeX aims to highlight these security risks to promote the safe use of DL compilers.

### v2

@@ III-D Model Reconstruction
For the global_scale method, different operators have same quant_scale, so the updated weights are close to the original model’s weights. For the kl_divergence method, quant_scale adjusts for different operators. Due to compilation optimizations, different quant_scale can be confused with regular multiplications, making it difficult to identify them. Eventually, the updated weights are a multiple of the actual weights, and different layers have different multiples. To address this, NeuroDeX adopts a substitute training method illustrated in Figure 5. The target model is the quantized compiled DNN executable, NeuroDeX first prepares query dataset to collect labels in the target executable. Next, NeuroDeX freezes known weights and trains the multiples between updated weights and actual weights. Typical black-box substitute method for model extraction (a in Figure 5) requires retraining all the model weights, whereas NeuroDeX only needs to train a limited multiples (b in Figure 5). A small amount of training data is enough to gain a model that effectively replicates the functionality of the target model.
@@ V-A RQ1: Correctness
TABLE V: TRA and THA of BTD (consistent with Table IV) on Different Compiler Versions (all value in %)
Metric TVM_O0 TVM_O3 GLOW
v0.7 v0.8 v0.9dev v0.7 v0.8 v0.9dev 2020 2021 2022
TRA_Avg 80.4 64.43 61.31 70.33 57.67 54.86 72.57 79.36 78.99
THA_Avg 96.42 94.57 98.50 98.04 97.05 97.10 97.33 98.48 96.91
We remove the training data that is also on test dataset (operators of the six models in Table IV) and retrain the BTD operator type recognition models following the default settings. The comparison results of TRA are shown in Table IV. The TRA of NeuroDeX significantly outperforms BTD, providing a crucial foundation for accurate operator type recognition, which is essential for subsequent operator attribute recovery and model reconstruction. NeuroDeX achieves results very close to 100% across all experimental models, whereas the BTD method faces relatively severe errors.
As a multi-classification machine learning model, BTD can also be evaluated based Type Hamming Accuracy (THA): $\frac{1}{N}\sum_{i=1}^{N}\left(\text{Pred}_{i}==\text{Label}_{i}\right)$ , where $N$ denotes the number of operator classes, and it’s also the length of the prediction vector for a single sample. The average result of TRA and THA for different compiler versions are shown in Table V. Although THA may achieve higher results, TRA undoubtedly provides a more scientific measure of the true effectiveness of operator recognition. For instance, prediction-label pair like $1,0,0,0,0\ldots]_{20}$ and $0,1,0,0,0\ldots]_{20}$ will yield THA with 18/20 = 0.9, but from the operator functional perspective, it is entirely incorrect.
@@ V-B RQ2: Efficiency
We choose four models: EfficientNet, ResNet18, Inceptionv1 and ShuffleNetv2. These models cover a range of weights size and topological complexities, enabling a comprehensive evaluation of the overhead associated with BTD and NeuroDeX. The model reconstruction strategies for NeuroDeX and BTD are identical. Therefore, we only compare the time associated with operator type recognition and operator attribute recovery processes. NeuroDeX involves interaction with LLM, which leads to non-fixed time consumption. We measure the time overhead by recording it three times and calculating the average to ensure a fair comparison. BTD uses IDA [37] to decompile the executables and NeuroDeX uses Ghidra. We overlook the difference from general decompiler preprocessing for fair comparison. The compiler version for our experiment is TVM v0.9dev and GLOW 2022. The overhead of BTD and NeuroDeX is shown in Table VI.
TABLE VI: Overhead Analysis on BTD and NeuroDeX (all value in s)
@@ V-C RQ3: Comprehensiveness
We first evaluate NeuroDeX on the latest compiler verison to verify its scalability. According to our observations, TVM is a project that is frequently maintained; GLOW is a stable project, we have counted the commits since 2023, which total only about 100, and the majority are related to feature maintenance and bug fixes. The repository of GLOW has been marked as public archive now. Therefore, we also use NeuroDeX to analyze the latest TVM version (v0.17). We utilize NeuroDeX to decompile RQ1’s six models in TVM 0.17 repeatedly. We evaluate NeuroDeX on six more different models in TVM v0.17, aiming to cover more model structures and varying hyperparameters for same model structure. The results are shown in Table VII.
TABLE VII: Evaluation of NeuroDeX on the Recent Compiler Version (all value in %)
@@
TABLE VIII: Evaluation of NeuroDeX on AArch64 Models (all value in %)
@@
In order to show the cross-architecture support of NeuroDeX, we also evaluate NeuroDeX on aarch64 in TVM v0.17, the results are shown in Table VIII. Based on the experimental results, NeuroDeX can successfully decompile executables across different optimization levels on aarch64 architecture. Some errors may accur in operator type recognition stage. The errors include the incorrect identification of Avgpool as Sum, as well as misjudgments of activation functions.
@@
For global_scale method, NeuroDeX can decompile all the four executables with MIA_top1 at least 26.4%, and MIA_top5 at least 40.6%. For kl_divergence method, NeuroDeX can decompile all the four executables with MIA_top1 at least 73.0%, and MIA_top5 at least 96.4%. Given that all models have 1000 output categories, the recovery results are quite satisfactory. In contrast, using the same computational graph with target model and random training with 5000 images (same query dataset with NeuroDeX), the recovered model achieves a maximum MIA_top1 of only 0.6% and a maximum MIA_top5 of only 4.4%, indicating that it almost fails to recover any valuable information. For ResNet18 and ResNet34, NeuroDeX can recover quite similar high-level models under both global_scale and kl_divergence methods. For MnasNet and MobileNetv2, the accuracy of recovered models decreases, we analyze the reasons as follows: compared to ResNet18 and ResNet34, MnasNet and MobileNetv2 have fewer parameters, which makes them more sensitive and difficult to determine the value of global_scale. The global_scale method tends to weaken the decision boundaries of models and cause the fluctuations in the model’s confidence. The queries for recovered model are more likely to be classified into different categories than original high-level models. This is due to the characteristics of the global_scale method and the inevitable precision loss during conversion between int and float domains, rather than design flaws in NeuroDeX.
@@
TABLE IX: TRA of Different LLMs (all value in %)
@@
V-D RQ4: Robustness
In operator type recognition, NeuroDeX may occasionally encounter errors. Additionally, due to the inherent output variability of LLMs, repeated analysis might yield different error samples. Nonetheless, our evaluation indicates that the operator type recognition accuracy on all TVM operators reach 99.22%, and it is 97.62% for GLOW, maintaining high accuracy. In operator attribute recovery, the errors are very rare and they are obvious in the decompiled code, such as kernel_size “*0.00591716 (1/169)” should be 13 but is incorrectly identified as 17. Therefore, it does not require a systematic error fix mechanism. We analyze the causes of errors in operator type recognition and the errors can be classified into three categories. Type1: a single operator is split into multiple functions; Type2: incorrect recognition of activation functions; Type3: other random operator recognition errors.
Type 1: Error operators exhibit fixed patterns. These errors accur in Softmax, Avgpool and Dense_add. These errors can be detected immediately after operator type recognition. NeuroDeX can fix them by matching predefined patterns of operator split.
To demonstrate the effectiveness of NeuroDeX on different LLMs, we also compare the performance of different LLMs including GPT-4.1 [38], Deepseekv3 [39], GPT-4o mini [40], Gemini2.5 flash [41]. The selection of LLMs will affect TRA directly. Results are shown in Table IX. Deepseekv3 and GPT-4.1 perform well in all models across different compiler settings. Gemini2.5 flash struggles to identify operators of GLOW, GPT-4o mini fails to identify TVM optimization operators and GLOW operators. In summary, NeuroDeX does not rely on a specific LLM. LLMs that perform well in general domains are equally suitable for completing the operator type recognition task.
@@ VI-A Possible Defenses towards NeuroDeX
NeuroDeX targets standard executables compiled using DL compilers and is capable of recovering nearly identical high-level models. Techniques such as binary obfuscation significantly reduce the readability of decompiled code by general-purpose decompilers like Ghidra. This can make NeuroDeX’s operator type recognition and operator attribute recovery components ineffective. However, DL compilers do not offer direct support for executable obfuscation and standardized solutions for obfuscating DNN executables have not been established. Methods like control flow obfuscation may also negatively impact the speed and accuracy of model inference. Overall, NeuroDeX is designed to handle the most common scenarios effectively. The threat model used in our study is consistent with previous works [9, 8, 7, 10] and is generally common and practical in real-world scenarios.
@@ VI-B Scalability of NLP Models
Related studies [42, 43] indicate that CV models are the majority models on edge devices, and our evaluation focuses on CV models. In contrast, language models have larger search spaces of input and output, as well as greater parameters. To the best of our knowledge, there is currently no mature solution for compiling them into standalone executables in binary format for inference on edge devices. Existing inference engines of LLMs primarily focus on optimizing the computation process but model weights are typically stored and loaded separately. Transformer-based models feature complex operator topologies, and operations like attention are decomposed into several low-level operators, further increasing the complexity of the models. This complexity makes it exceedingly difficult to precisely reconstruct an identical model. NeuroDeX’s core components are still effective for NLP models, but adapting them for NLP models requires a significant amount of work including heuristic methods, we will address this in the future work.